# Revisiting Pre-trained Language Models for Vulnerability Detection

### v1

@@ II-B Evaluating PLMs for Vulnerability Detection
Limitations in Existing Evaluations. Existing works fall short in comprehensiveness and accuracy of evaluation, limiting our understanding of LLMs’ applicability in real-world VD applications. These limitations are mainly reflected in the limited scope of models, datasets, and adaptation methods chosen for evaluation, as well as in flawed data partitioning methods and experimental settings, and a lack of analysis of models’ generalization and robustness in real-world scenarios.
@@
First, many prior studies focus exclusively on either fine-tuning [1, 2, 3, 4, 5, 6, 7, 8] or prompt engineering [32, 19, 20, 25, 26, 27, 28, 21, 22, 23, 24, 18], without conducting a comprehensive comparison or analyzing the trade-offs between performance and efficiency. Also, some evaluations suffer from biased evaluation results due to limitations in model size [18, 1, 2, 3, 5, 6, 7, 9, 11, 13, 14], dataset scale, and vulnerability diversity [19, 20, 25, 27, 21, 22, 23, 24, 2, 5, 7, 11, 12, 13]. Furthermore, inappropriate data partitioning can skew training outcomes and produce unrealistic performance estimates. In real-world settings, VD datasets are often highly imbalanced. Fine-tuning under such conditions can bias models toward non-vulnerable samples, leading to high false negative rates [1, 9, 33, 10, 18]. Conversely, evaluating models under artificially balanced conditions risks grossly overestimating precision [32, 19, 20, 25, 26, 27, 28, 21, 22, 23, 24, 2, 3, 4, 17, 11, 13], as false positives are just as critical in security analysis as false negatives. Another concern is the prevalent use of random data partitioning [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], which risks data leakage if similar vulnerability patterns appear in both training and test sets. Potential temporal overlap between a model’s pretraining data cutoff and the test set’s commit dates can further exacerbate this issue [32, 19, 25, 28, 21, 22, 24, 18, 1, 2, 3, 4, 5, 6, 7, 8, 9, 33, 10, 16, 17, 11, 12, 13, 14, 15]. Moreover, evaluations conducted solely on in-distribution data may overestimate model performance [2, 6, 8, 9, 10, 16, 17, 11, 12, 13, 14]. Validating models on test data from diverse sources is essential to assess generalization. Finally, robustness analysis is equally important, as it evaluates how models withstand real-world adversarial challenges, such as detecting vulnerable code reuse under various perturbations.
@@
Our Work. In contrast to limitations of existing evaluations as detailed in Table I, our study comprehensively addresses all of aforementioned issues at every stage of the evaluation process. For the first time, we conduct an in-depth and accurate evaluation of 17 representative PLMs for VD, encompassing a variety of model architectures and systematically compares fine-tuning and prompt engineering on a self-collected VD dataset. For fine-tuning, we include open-source PLMs with up to 34 billion parameters. For prompt engineering, we evaluate two prompt settings and four types of structure- and semantic-aware prompts to determine optimal configurations. Finally, we assess the real-world performance of PLMs from multiple realistic perspectives, including their generalization to out-of-distribution data and robustness to code perturbations (normalization, abstraction, semantic-preserving transformations, truncation). Our analysis offers valuable insights to guide future research in VD using PLMs.
III Experimental Setup
III-A Reconstructed Dataset
III-A1 Dataset Selection
@@
Despite efforts to build numerous VD datasets and benchmarks, limitations like limited data diversity/volume, unrealistic synthetic samples, balanced evaluation data distributions, and inaccurate labeling still persist, leading to biased evaluations in VD research. Many existing datasets cover only a few CWE types or projects [45, 46, 47, 48], or include synthetic data that poorly reflects real-world vulnerabilities [49, 50, 51, 52]. While some datasets like SVEN [53] and SecLLMHolmes [20] are collected from real-world projects and manually labeled, they are limited in scale and exhibit unrealistic class balance. More comprehensive datasets like Big-Vul [54], CVEFixes [55], PatchDB [56], CrossVul [57], DiverseVul [1], and MegaVul [58] span multiple programming languages or CWE types. However, their labeling approach, which tags pre-patch functions as vulnerable, can result in false positives by mislabeling unrelated code changes. To address these issues, PrimeVul [33] combines four major VD datasets and applies refined labeling rules, improving labeling accuracy by 26%–67%. Given its enhanced diversity and accuracy, we construct our evaluation dataset based on PrimeVul.
III-A2 Dataset Partitioning
@@
In real-world applications, VD models should learn from historical vulnerabilities to detect future ones. However, most prior studies [10, 1, 9, 47, 59] randomly split datasets (8:1:1 ratio), risking data leakage. For instance, a patched (non-vulnerable) function may appear in training, while its pre-patch (vulnerable) version is in testing (Case 1). Additionally, similar vulnerability fixes within the same commit (e.g., QEMU 902b27d [60]) may be distributed across different data splits, potentially causing the similar patterns learned during training to reappear in the evaluation set (Case 2). To mitigate this, it is necessary to introduce a time-based partitioning method based on commit date [33].
@@
Also, random splitting could lead to the distribution of different vulnerability types being misaligned with real-world scenarios, resulting in biased conclusions (e.g., Table 7 in DiverseVul [1]). For instance, some common CWE types may be underrepresented in training set, which hinders the model from learning relevant vulnerability patterns. Conversely, certain rare CWE types might be mostly allocated to the training set, producing evaluation results on the sparse test data that do not accurately reflect the model’s real-world performance.
@@
III-B Self-Collected Dataset
Due to potential overlap between the evaluation data for VD and the pre-training data for PLMs (e.g., both collected from same GitHub commits) [61], some studies have independently curated holdout test sets to avoid data leakage and minimize bias in evaluation. However, these datasets often exhibit limitations in terms of data scales, project diversity, and coverage of CWE types. For instance, Ullah et al. [20] curated a dataset consisting of only 228 samples, covering 8 CWE types from 4 projects, which may limit the generalizability of their findings.
To this end, we curate a new VD dataset by collecting all C/C++ vulnerability-related GitHub commits from NVD with commit dates spanning the period from October 2023 to October 2024, which postdate the pre-training cutoff dates of PLMs evaluated in this study. For each commit, we record detailed metadata about the vulnerability and project information. We segment functions and label them following the prior work [54, 33]: functions before patching are labeled as vulnerable, while functions after patching and unchanged functions are labeled as non-vulnerable. Ultimately, we obtain a evaluation dataset comprising 25,536 functions, including 646 vulnerable functions and 24,890 non-vulnerable functions, spanning 99 projects and 28 CWE types.
Note that our automated process used during dataset construction can be easily extended to collect and process the latest vulnerability/patch commits from NVD. We will release the artifacts to facilitate future research and enable seamless expansion of our collected dataset for research purposes.
III-C Statistics of the datasets used in evaluations
To investigate the performance of PLMs under different training/test settings (Section IV-A), we construct a balanced training set based on our reconstructed training set, which by default is imbalanced. Specifically, we adjust the ratio of vulnerable to non-vulnerable functions in the training set to 1:1 through random undersampling, while keeping the original validation and test sets unchanged. The self-collected dataset serves as an out-of-distribution test set to evaluate generalizability while avoiding data leakage (Section IV-B and IV-A). The statistics for the datasets used in our evaluation are provided in Table II.
TABLE II: Statistics of the datasets used in evaluations
@@ III-C Statistics of the datasets used in evaluations
III-D Evaluated Models
To ensure comprehensive evaluations, we select a set of representative PLMs covering encoder-based (CodeBERT, GraphCodeBERT, UniXCoder, PDBERT), decoder-based (CodeLlama-7B, 13B and 34B, DeepSeek-Coder-6.7B and 33B, StarChat-$\beta$-16B, WizardCoder-15B and 33B, Mistral-7B, GPT-3.5, GPT-4, GPT-4o Mini), and encoder-decoder based model architectures (CodeT5). The rationale for selecting the above PLMs is that all of them are oriented toward code-specific tasks or have been fed on a substantial amount of code data during pre-training. This endows them with a foundational understanding of code, which proves advantageous when applying them to downstream VD tasks.
III-E Prompt Design
Prompt Settings. We evaluate the performance of LLMs in VD under two prompt settings: (1) zero-shot prompting: directly asking the LLM if the function is vulnerable without providing any additional information; (2) few-shot in-context learning (ICL): providing the LLM with 4 shots (2 randomly selected patched pairs each time) beforehand, with each shot including a query and its ground truth answer, followed by asking the LLM whether a new function is vulnerable.
Under each prompt setting, we categorize the prompts into four types: ① raw code, ② flattened AST, ③ code with API calls, and ④ code with data flow. In ②, ③, and ④, the prompt contains not only raw code but also structural information within the code, to explore if this assists the LLM in understanding the code and its vulnerability patterns. We use Tree-sitter [62] to extract the structural information from the code and serialize it into plain text. Note that ② and ④ are prompt types newly proposed and evaluated in this work.
For ②, after obtaining the AST of code, we adopt UniXCoder [37]’s AST mapping algorithm to generate flattened AST. For example, the code line c=a+b; is translated to:
[…]
@@ III-E Prompt Design
[…]
For ③, we traverse the AST and collect nodes of type “call_expression”. Following prior work [18], we describe the flow of API call using the following template:
[…]
For ④, inspired by GraphCodeBERT [38], which incorporates data flow during pre-training to enhance code understanding, we explore the effect of introducing data flow descriptions during inference. We extract data flow between variable nodes in the format (VAR, $p_{i}$, comesFrom, [VAR], [$p_{j}$]), indicating that the variable VAR at position $p_{i}$ originates from another at $p_{j}$. While [18] describes this as “the data value of VAR at the $p_{i}$th token comes from data at the $p_{j}$th token”, using absolute token positions could misalign with LLM tokenization, leading to incorrect interpretation. To ensure that the LLM accurately identifies tokens and understands the data flow, we use the relative position of the node VAR within the function and translate the above example as below:
[…]
Prompt Templates. To guide LLMs in making better predictions, our prompts are composed of two components: system role and user content. For the system role, the prompts begin with the instruction: “You are a code security expert who excels at detecting vulnerabilities”. The user content starts with the question: “Is the following function vulnerable? Please answer Yes or No”, followed by the input function. Additionally, we adopt the default chat templates of LLMs by calling tokenizer.apply_chat_template() [63] to align with their instruction data format.
TABLE III: Comparison of PLMs under imbalanced and balanced training settings
@@ III-F Metrics
III-G Implementation
III-G1 Evaluation Framework.
@@ III-G Implementation
Given the variations in implementation across existing empirical studies on PLMs for VD [10, 33, 7, 19, 20, 30], we develop a comprehensive evaluation framework for PLMs in VD, building upon  [33, 19]. This framework is designed to be easily extensible to new datasets, PLMs, model adaptation techniques, with the goal of enhancing the accuracy and fairness of PLM evaluations for VD. Specifically, for open-source PLMs, we utilize the Hugging Face transformers library [64] to load configurations, tokenizers, and base models. For SLMs, we perform full fine-tuning. For LLMs, we use LoRA to fine-tune the models and incorporate DeepSpeed ZeRO-3 along with the Accelerate module to improve training efficiency. For GPT models, we use the Azure OpenAI API for inference.
@@
All experiments are conducted on a computational node equipped with multiple NVIDIA A100 GPUs (80 GB VRAM each), 64 CPU cores, and 512 GB of RAM. For GPT-related inference, approximately 38 million tokens are processed per model, with a total cost of \$2,400.
III-G2 Hyperparameters

### v3

@@ 2.3. Evaluating PLMs for Vulnerability Detection
Limitations in Existing Evaluations. Existing works fall short in comprehensive and accurate evaluations, which limits our understanding of LLMs’ applicability in real-world VD applications. These limitations stem from a limited scope, resulting from evaluations of constrained evaluation setups on unrepresentative experimental settings. They also arise from flawed data partitioning methods, overlooked temporal overlap between evaluation data and LLM pre-training knowledge, and a lack of analysis regarding model generalization and robustness in real-world scenarios.
@@
First, many prior studies focus exclusively on either fine-tuning (diversevul; a16; a17; a18; a19; a20; benjamin2023; a22) or prompt engineering (understanding; benjamin; ullah; a6; a7; a8; a9; a10; a11; a12; a13; graph_prompt), without conducting a comprehensive comparison or analyzing the trade-offs between performance and efficiency. Also, some evaluations suffer from biased evaluation results due to limitations in model size (graph_prompt; diversevul; a16; a17; a19; a20; benjamin2023; ni1; a28; davidlo; a31), dataset scale, and vulnerability diversity (benjamin; ullah; a6; a8; a10; a11; a12; a13; a16; a19; benjamin2023; a28; a29; davidlo). Furthermore, inappropriate data partitioning can skew training outcomes and produce unrealistic performance estimates. In real-world settings, VD datasets are often highly imbalanced. Fine-tuning under such conditions (diversevul; ni1; primevul; yinxin; graph_prompt) can bias models toward non-vulnerable samples, leading to high false negative rates. Conversely, evaluating models under artificially balanced conditions (understanding; benjamin; ullah; a6; a7; a8; a9; a10; a11; a12; a13; a16; a17; a18; a27; a28; davidlo) risks grossly overestimating precision, as false positives are just as critical in security analysis as false negatives. Another concern is the prevalent use of random data partitioning (diversevul; a16; a17; a18; a19; a20; benjamin2023; a22; ni1; yinxin; a28; a29; davidlo; a31; a32), which risks data leakage if similar vulnerability patterns appear in both training and test sets. Potential temporal overlap  (understanding; benjamin; a6; a9; a10; a11; a13; graph_prompt; diversevul; a16; a17; a18; a19; a20; benjamin2023; a22; ni1; primevul; yinxin; a26; a27; a28; a29; davidlo; a31; a32) between LLMs’ pretraining data cutoff dates and test data’s commit dates can further exacerbate this issue. Moreover, evaluations conducted solely on in-distribution data (a16; a20; a22; ni1; yinxin; a26; a27; a28; a29; davidlo; a31) may overestimate model performance. Validating models on test data from diverse sources is essential to assess generalization. Finally, robustness analysis is equally important, as it evaluates how models withstand real-world adversarial challenges, such as detecting vulnerable code reuse under various perturbations.
@@
Our Work. In contrast to limitations of existing evaluations as detailed in Table 1, our study comprehensively addresses all of aforementioned issues at every stage through the evaluation pipeline. For the first time, we conduct the in-depth and accurate evaluation of 18 representative PLMs for VD, encompassing different scales of model parameters and systematically compares fine-tuning and prompt engineering on the high-quality datasets. For fine-tuning, we include open-source PLMs with up to 34 billion parameters. For prompt engineering, we evaluate two prompt settings and four types of structure- and semantic-aware prompts to determine better configurations. Finally, we assess the real-world performance of PLMs from multiple realistic perspectives, including their generalization to out-of-distribution data and robustness to code perturbations (normalization, abstraction, semantic-preserving transformations, truncation). Our analysis offers valuable insights to guide future research in VD using PLMs.
3. Experimental Setup
3.1. Reconstructed Dataset
3.1.1. Dataset Selection
@@
Despite efforts to build numerous VD datasets and benchmarks (bigvul; crossvul; diversevul; Devign; ReVeal), limitations like limited data diversity/volume, unrealistic synthetic samples, balanced evaluation data distributions, and inaccurate labeling still persist, leading to biased evaluations in VD research. Many existing datasets cover only a few CWE types or projects (VulDeePecker; Devign; ReVeal; DeepWukong), or include synthetic data that poorly reflects real-world vulnerabilities (uVulDeePecker; juliet:online; SySeVR; VulDeeLocator). While some datasets like SVEN (SVEN) and SecLLMHolmes (ullah) are collected from real-world projects and manually labeled, they are limited in scale and exhibit unrealistic class balance. More comprehensive datasets like Big-Vul(bigvul), CVEFixes (cvefixes), PatchDB (wang2021patchdb), CrossVul (crossvul), DiverseVul (diversevul), and MegaVul (megavul) span multiple programming languages or CWE types. However, their labeling approach, which tags pre-patch functions as vulnerable, can result in false positives by mislabeling unrelated code changes. To address these issues, PrimeVul (primevul) combines four major VD datasets and applies refined labeling rules, improving labeling accuracy by 26%–67%. Given its enhanced diversity and accuracy, we construct our evaluation dataset based on PrimeVul.
3.1.2. Dataset Partitioning
@@
In real-world applications, VD models should learn from historical vulnerability-related samples to detect future ones. However, most prior studies (yinxin; diversevul; ni1; ReVeal; linevul) randomly split datasets (8:1:1 ratio), risking data leakage. For instance, a patched (non-vulnerable) function may appear in training, while its pre-patch (vulnerable) version is in testing (Case 1). Additionally, similar vulnerability fixes within the same commit (e.g., QEMU 902b27d (QEMU_commit)) may be distributed across different data splits, potentially causing the similar patterns learned during training to reappear in the evaluation set (Case 2). To mitigate this, it is necessary to split the dataset based on commit date (primevul).
@@
Furthermore, random splitting could lead to the distribution of different vulnerability types being misaligned with real-world scenarios, resulting in biased conclusions (e.g., Table 7 in DiverseVul (diversevul)). For instance, some common CWE types may be underrepresented in the training set, which hinders the model from learning relevant vulnerability patterns. Conversely, certain rare CWE types might be mostly allocated to the training set, producing evaluation results on the sparse test data that do not accurately reflect the model’s real-world performance.
@@
3.2. Self-Collected Dataset
Due to potential overlap between the evaluation data for VD and the pre-training data for PLMs (e.g., both collected from same GitHub commits) (quality), some studies have independently curated holdout test sets to avoid data leakage and minimize bias in evaluation. However, these datasets often exhibit limitations in terms of data scales, project diversity, and coverage of CWE types. For instance, Ullah et al. (ullah) curated a dataset consisting of only 228 samples, covering 8 CWE types from 4 projects, which may limit the generalizability of their findings.
To this end, we curate a new VD evaluation dataset by collecting all C/C++ vulnerability-related GitHub commits from NVD with commit dates spanning the period from October 2023 to October 2024, which postdate the pre-training cutoff dates of PLMs evaluated in this study. For each commit, we record detailed metadata about the vulnerability and project information. We segment functions and label them following the prior work (bigvul; primevul): changed functions before patching are labeled as vulnerable, while changed functions after patching and unchanged functions are labeled as non-vulnerable. Ultimately, we obtain a evaluation dataset comprising 25,536 functions, including 646 vulnerable functions and 24,890 non-vulnerable functions, spanning 99 projects and 28 CWE types. Fig. 1 shows the distribution of top 10 vulnerability types of data in our self-collected dataset. Over half of the CWE types are listed among the 2024 top 25 most dangerous software weaknesses (cwe2024top25), demonstrating the comprehensiveness of our evaluated data. Note that our automated dataset construction can be easily extended to collect and process the latest patch commits from NVD.
Figure 1. Distribution of top 10 CWEs in self-collected data:
3.3. Statistics of the Datasets Used in Evaluations
Table 2. Statistics of the datasets used in evaluations
@@ 3.3. Statistics of the Datasets Used in Evaluations
To investigate the performance of PLMs under different training and test settings (Section 4.1), we construct a balanced training set based on our reconstructed training set, which by default is imbalanced. Specifically, we adjust the ratio of vulnerable to non-vulnerable functions in the training set to 1:1 through random undersampling, while keeping the original validation and test sets unchanged. The self-collected dataset serves as an out-of-distribution test set to evaluate generalizability while avoiding data leakage (Section 4.1 and 4.2). The statistics for the datasets used in our evaluation are provided in Table 2.
3.4. Evaluated Models
To ensure comprehensive evaluations, we select a set of representative PLMs covering encoder-based (CodeBERT, GraphCodeBERT, UniXCoder, PDBERT), decoder-based (CodeLlama-7B, 13B and 34B, DeepSeek-Coder-6.7B and 33B, StarChat-$\beta$-16B, WizardCoder-15B and 33B, Mistral-7B, GPT-3.5, GPT-4, GPT-4o Mini, and GPT-5 Mini), and encoder-decoder based model architectures (Code T5). Table 3 and Appendix B show the details of models evaluated in this work.
The rationale for selecting the above PLMs is that all of them are oriented toward code-specific tasks or have been fed on a substantial amount of code data during pre-training. This endows them with a foundational understanding of code, which proves advantageous when applying them to downstream VD tasks.
Table 3. Models evaluated in this work
Model
Category
Model
API
Parameter
Size
Max Position
Embeddings
Knowledge
Cutoff Date
SLMs
codebert-base
graphcodebert-base
unixcoder-base-nine
codet5-base
pdbert-base
125M
125M
125M
220M
125M
512
512
512
512
512
<Jul 2020
<Feb 2021
<Aug 2021
<Sep 2021
<Nov 2023
Open-source
LLMs
deepseek-coder-6.7b-instruct
deepseek-coder-33b-instruct
Mistral-7B-Instruct-v0.1
CodeLlama-7b-Instruct-hf
CodeLlama-13b-Instruct-hf
CodeLlama-34b-Instruct-hf
WizardCoder-15B-V1.0
WizardCoder-33B-V1.1
starchat-beta
6.7B
33B
7B
7B
13B
34B
15B
33B
16B
16K
16K
32K
16K
16K
16K
8K
16K
8K
Mar 2023
Mar 2023
Aug 2021
Jan 2023
Jan 2023
Jan 2023
<Jun 2023
<Jun 2023
<Jun 2023
Closed-source
LLMs
gpt-3.5-turbo-0125
gpt-4-32k
gpt-4o-mini
gpt-5-mini
7B
16K
32K
128K
400K
Sep 2021
Sep 2021
Oct 2023
Jun 2024
3.5. Prompt Design
Prompt Settings. We evaluate the performance of LLMs in VD under two prompt settings: (1) zero-shot prompting: directly asking the LLM if the function is vulnerable without providing any additional information; (2) few-shot in-context learning (ICL): providing the LLM with 4 shots (2 randomly selected pre-patched and post-patched function pairs each time) beforehand, with each shot including a query and its ground truth answer, followed by asking the LLM whether a target function is vulnerable.
Under each prompt setting, we categorize the prompts into four types: ① raw code, ② flattened AST, ③ code with API calls, and ④ code with data flow. In ②-④, the prompt contains not only raw code but also structural information within the code, to explore if this assists the LLM in understanding the code and its vulnerability patterns. We use Tree-sitter (treesitter) to extract the structural information from the code and serialize it into plain text.
For ②, after obtaining the AST of code, we adopt UniXCoder (UniXcoder)’s AST mapping algorithm to generate flattened AST. For example, the code line c=a+b; is translated to:
@@ 3.5. Prompt Design
<AST#expression_statement#Right>
For ③, we traverse the AST and collect nodes of type “callexpression”. Following prior work (graph_prompt), we describe the flow of API call using the following template:
API call: The program first calls <node_0>, … , then calls <node_1>, … , and finally calls <node_n>.
For ④, inspired by GraphCodeBERT (graphcodebert), which incorporates data flow during pre-training to enhance code understanding, we explore the effect of introducing data flow descriptions during inference. We extract data flow between variable nodes in the format (VAR, $p_{i}$, comesFrom, [VAR], $p_{j}$]), indicating that the variable VAR at position $p_{i}$ originates from another at $p_{j}$. While (graph_prompt) describes this as “the data value of VAR at the $p_{i}$th token comes from data at the $p_{j}$th token”, using absolute token positions could misalign with LLM tokenization, leading to incorrect interpretation. To ensure that the LLM accurately identifies tokens and understands the data flow, we use the relative position of the node VAR within the function and translate the above example as below:
Data flow: The 2nd VAR comes from the 1st VAR …
Prompt Templates. To guide LLMs in making better predictions, our prompts are composed of two components: system role and user content. For the system role, the prompts begin with the instruction: “You are a code security expert who excels at detecting vulnerabilities”. The user content starts with the question: “Is the following function vulnerable? Please answer Yes or No”, followed by the input function. Additionally, we adopt the default chat templates of LLMs by calling tokenizer.apply_chat_template() (huggingface_chat_templates) to align with their instruction data format.
Table 4. Comparison of PLMs under imbalanced and balanced training settings
@@
[…]
14.68 | 12.77
14.15 | 11.87
13.33 | 10.49
17.18 | 14.18
13.98 | 11.72
8.19 | 7.30
7.79 | 6.60
7.29 | 5.74
9.62 | 8.03
7.76 | 6.53
70.75 | 50.77
77.67 | 58.82
77.38 | 61.46
79.83 | 60.53
70.61 | 56.81
75.48 | 83.27
71.57 | 78.40
69.59 | 73.79
76.82 | 82.01
74.04 | 78.90
3.6. Metrics
We consider multiple metrics to comprehensively demonstrate the models’ capabilities, formulated as follows: Accuracy = $(TP+TN)/(TP+TN+FP+FN)$, Recall = $TP/(TP+FN)$, Precision = TPTP/$(TP+FP)$, True Negative Rate (TNR) = $TN/(TN+FP)$, F1 = $2\cdot(\text{Precision}\cdot\text{Recall})/(\text{Precision}+\text{Recall})$, Balanced Accuracy = $(\text{Recall}+\text{TNR}))/2$. Since we expect VD models to exhibit both low false positive and false negative rates in practice, following previous work (benjamin), we use balanced accuracy as our main metric to balance true positive predictions and true negative predictions.
3.7. Implementation
3.7.1. Evaluation Framework
@@
Given the variations in implementation across existing empirical studies on PLMs for VD (yinxin; primevul; benjamin2023; benjamin; ullah; pdbert), we develop a comprehensive evaluation framework for PLMs in VD, building upon  (primevul; benjamin). This framework is designed to be easily extensible to new datasets, PLMs, model adaptation techniques, with the goal of enhancing the accuracy and fairness of PLM evaluations for VD. Specifically, for open-source PLMs, we utilize the Hugging Face transformers library (hf) to load configurations, tokenizers, and base models. For SLMs, we perform full fine-tuning. For LLMs, we use LoRA to fine-tune the models and incorporate DeepSpeed ZeRO-3 along with the Accelerate module to improve training efficiency. For GPT models, we use the Azure OpenAI API for inference.
@@
All experiments are conducted on a computational node equipped with multiple NVIDIA A100 GPUs (80 GB VRAM each), 64 CPU cores, and 512 GB of RAM. For GPTs, approximately 38 million tokens are processed per model, with a total cost of \$2,400.
3.7.2. Hyperparameters